Suppose you deposited $3,000 in a savings account earning 3.1% interest compounding daily. How long will it take for the balance to grow to $7,000? Answer in years rounded to one decimal place.
Answer #1

We use the formula:
A=P(1+r/365)^365n
where
A=future value
P=present value
r=rate of interest
n=time period.

7000=3000*(1+0.031/365)^365n

(7000/3000)=(1.000084932)^365n

Taking log on both sides;

log (7000/3000)=365n*log(1.000084932)

n=1/365[log (7000/3000)/log(1.000084932)]

=27.3 years(Approx).